payton:

is 5(x-2)=5x-7 a... one solution no solution infinitely many sloutions

TheSmartOne:

First distribute the 5 What do you get when you distribute the 5 5(x-2)

payton:

5x and -10

TheSmartOne:

Perfect so now we have 5x - 10 = 5x - 7 Subtract 5x on both sides, what do you get?

payton:

-10=x-7

TheSmartOne:

mhmm I'm not so sure about that 5x - 10 - 5x = 5x - 7 - 5x Simplify that Remember, we can only subtract the like terms so whatever has an x with it

payton:

wait i dont get it

TheSmartOne:

think of 5x as five apples we have Five apples - 10 = five apples - 7 When we do one thing to one side, we have to do it to the other side so that way it's balanced So like for example If we have 3 = 3 That's true right? Three equals to three If I add 1 on only one side then it becomes 4 = 3 Then it doesn't stay true anymore That's why I have to add one to both sides And we get 4= 4 And that still keeps it true!

TheSmartOne:

Do you get that now

payton:

so im going to subtract 5 by -10 and by -7?

TheSmartOne:

No no no Like terms So we have 5x - 10 The five has an x The ten has nothing with it One is an apple the other is an orange You can't add or subtract them When the number has a letter next to it, you can't add it or subtract it with anything except another number with the SAME letter So 5x + 10x = 15x 5x + 2y = 5x + 2y CAN'T ADD 2a + 3a = 5a 3b + 5 = 3b + 5 CAN'T ADD

TheSmartOne:

Does that make sense to you?

payton:

yes

payton:

oh wait is it -10 = -7

payton:

so that would be no solution?

TheSmartOne:

Exactly!! You got it -10 can never equal to -7 So it's no solutions

payton:

okay